OpenStudy (anonymous):

Limit Problem! What is the limit of, as x approach infinity, of (x^3)*e^(-x^2)? If someone could help me answer this problem, I'd greatly appreciate it! Thanks!

OpenStudy (anonymous):

you know Lhopital's Rule?

OpenStudy (anonymous):

mhmm, that's what i've been trying to use

OpenStudy (anonymous):

you'll prolly have to apply it more than once.

OpenStudy (anonymous):

I've been trying it. But I can't arrive at an answer.

OpenStudy (anonymous):

how many times did you apply it?

OpenStudy (anonymous):

I just looked. I'm pretty sure twice.

OpenStudy (anonymous):

Looking over at my computation, I think the answer is infinity. Can you verify?

OpenStudy (anonymous):

you can write the expression as \[\large \frac{x^3}{e^{x^2}} \]

OpenStudy (anonymous):

Oh yeah, you're right! I didn't think of it that way. Lemme try apply the rule once more after the derivative.

OpenStudy (anonymous):

the denominator will still have that e^(x^2) the numerator will end up a constant.... shouldn't the limit be 0 then?

OpenStudy (anonymous):

I'll will try and compute first using the quotient rule, correct?

OpenStudy (anonymous):

no... quotient rule is not Lhopital's Rule.

OpenStudy (anonymous):

just take derivative of the top then derivative of the bottom

OpenStudy (anonymous):

But don't I need to differentiate first?

OpenStudy (anonymous):

Since it's a quotient, shouldn't I use the quotient rule though?

OpenStudy (anonymous):

\[\large lim \frac{f}{g} =lim\frac{f'}{g'}\] this is Lhopitals Rule

OpenStudy (anonymous):

where f=x^3 and g=e^(x^2)

OpenStudy (anonymous):

uh-huh. then i got that to equal: \[3x ^{2}/e ^{x ^{2}}\]

OpenStudy (anonymous):

your denominator, g' = 2xe^(x^2)

OpenStudy (anonymous):

it goes to 0

OpenStudy (anonymous):

you can tell as you take more and more derivatives the top goes to 0 where the denominator never will

OpenStudy (anonymous):

then it looks like you'll have to apply it once more since you still have an indeterminate form after applying Lhopital's rule the first time.

OpenStudy (anonymous):

But then the 2x in the denominator can cancel out with the numerator. Wouldn't it be: 3x/2e^x^2

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

but you still have an indeterminate form here still...

OpenStudy (anonymous):

so .... once more....:)

OpenStudy (anonymous):

like a drill sargeant.

OpenStudy (anonymous):

Okay, I did it once more. And I got 3/4xe^x^2.

OpenStudy (anonymous):

so as x approaches infinity.... what does the valu of that expression approach?

OpenStudy (anonymous):

0! Gotcha. But is my last differentiation correct?

OpenStudy (anonymous):

yes it is...:)

OpenStudy (anonymous):

Ah, great! Thank you so much! That was not as bad as I first thought it would be. I just completely blanked out with the negative exponent. I forgot you could place it in the denominator. But I understand now. Thank you!

OpenStudy (anonymous):

yw...:)